OpenStudy (anonymous):

3. How many solutions does the system of equation have? 3x=-12y+ 15 and + 4y = 5 a. one b. two c. infinitely many d. none

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

3x=-12y+ 15 4y = 5 your two equations correct?

OpenStudy (anonymous):

3x= -12y+15 and x+4y=5*

OpenStudy (dtan5457):

Infinite solutions. isolate y in both equations 12y=15-3x 4y=5-x divide the first one by 3 and you get the same thing as the 2nd one. same line, infinite solutions

OpenStudy (dtan5457):

questions?

OpenStudy (anonymous):

oh wait!

OpenStudy (anonymous):

a. one b. two c. infinitely many d. none

OpenStudy (anonymous):

@dtan5457 I forgot to put the answer choices

OpenStudy (dtan5457):

well exactly i said infinite so c. infinitely many

OpenStudy (anonymous):

oooo okay! I got more questions \(^0^)/

OpenStudy (anonymous):

4. how many solutions does the system of equations have y= 6x + 2 y= 6x + 4 a. one b. two c. infinite d. none

OpenStudy (dtan5457):

okie. i want you to solve that one. what do you think.

OpenStudy (anonymous):

D: don't make meee D:

OpenStudy (dtan5457):

i can't save you during a real test lol

OpenStudy (dtan5457):

let me make this easier

OpenStudy (anonymous):

okii

OpenStudy (dtan5457):

I graphed both lines. from what i said yesterday, how many solutions does this graph show?

OpenStudy (dtan5457):

Do you see them? lol

OpenStudy (anonymous):

yes. and it's two lines

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

ok...but these lines are parallel.

OpenStudy (dtan5457):

what does that mean to you

OpenStudy (anonymous):

there not intersecting

OpenStudy (dtan5457):

so many solutions lol

OpenStudy (dtan5457):

*how many solutions

OpenStudy (anonymous):

none?

OpenStudy (dtan5457):

correct ^

OpenStudy (anonymous):

YAYYY

OpenStudy (dtan5457):

say NONE! not "none?" NO INTERSECTIONS=NO SOLUTION

OpenStudy (anonymous):

NONE!!!

OpenStudy (dtan5457):

please don't forget that lol

OpenStudy (anonymous):

I'll try. but I even had a dream last night that I was taking my algebra Exam and I freaked out

OpenStudy (anonymous):

I have more questionssss

OpenStudy (dtan5457):

lol relax. go ask more questions.

OpenStudy (anonymous):

x-\[x-3y=15 and y= \frac{ 1 }{ 3 }x - 5\] a. one b. two c. none d. infinitely many

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

to clarify your equations: x-3y=15 y=1/3x-5

OpenStudy (anonymous):

yess.

OpenStudy (dtan5457):

isolate y in the first equation please

OpenStudy (dtan5457):

x-3y=15

OpenStudy (dtan5457):

get y alone

OpenStudy (dtan5457):

try please

OpenStudy (anonymous):

o.o you want me to try? okay...

OpenStudy (dtan5457):

LOL yes. all your questions are pretty much the same. I'm gonna make you do this whole problem, it's really not that hard once you get the concepts

OpenStudy (dtan5457):

you don't have school tomorrow correct?

OpenStudy (anonymous):

correct but I've been doing this math literally all day and I'm almost doneee D:

OpenStudy (anonymous):

6 more questionsssss D:

OpenStudy (dtan5457):

did you do them without my help lol

OpenStudy (dtan5457):

if you knew how to do these problems you'll do your last 6 questions in under 10 minutes..

OpenStudy (dtan5457):

by yourself lol

OpenStudy (anonymous):

UGHHH HYKFJNKJNEKJ I did the others with desmo but then I realized I couldn't do these with desmos and now Im screweddd D:

OpenStudy (dtan5457):

actually you can do them with desmos as long as you isolate y

OpenStudy (dtan5457):

isolate y missy

OpenStudy (anonymous):

shhhhh let's act like I can't D:

OpenStudy (anonymous):

shhhh shhhh

OpenStudy (anonymous):

I hate math -.-

OpenStudy (anonymous):

I got a question...(im sorry im done and I forget things) but how do I get y alone again?

OpenStudy (dtan5457):

x-3y=15 put y on one side of the equal sign with the other terms to the other side. if your gonna remove the x you subtract it so don't forget the negative sign

OpenStudy (dtan5457):

oh yeah, and just a hint these are linear equations there will never be 2 solutions its either 1, none, or infinite

OpenStudy (dtan5457):

so right now let me see your equation after you isolated y please

OpenStudy (anonymous):

oohhh kkkk

OpenStudy (anonymous):

well *nervously laughs* about go isolate that y mhm okay

OpenStudy (dtan5457):

look if i told you 5+5=10 i can switch this equation however i want, i can subtract 5 on each side and still get 5=5

OpenStudy (anonymous):

oh because 10-5 is 5

OpenStudy (dtan5457):

YES

OpenStudy (dtan5457):

so for x-3y=15

OpenStudy (dtan5457):

in what way is y not alone

OpenStudy (dtan5457):

hint: there's a frickin x right in front of it

OpenStudy (dtan5457):

remove it mate

OpenStudy (anonymous):

wait so am I subtracting 15 on each sides? (sorry, stupid question OMG IM SO SORRY FOR BEING STUPID. LIKE OMG IM SORRY LIKE SORRY IM SORRY)

OpenStudy (anonymous):

*is a failure*

OpenStudy (dtan5457):

you could but it just makes it more confusing x-3y=15 remove the x SUBTRACT THE X

OpenStudy (anonymous):

how do you subtract x? like actually put x- 15??? or am I putting x-3?? AHHHHHHH

OpenStudy (dtan5457):

x-3y=15 -x -x =???

OpenStudy (anonymous):

*runs around in circles*

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

wait do I put this in desmo?

OpenStudy (anonymous):

o.o

OpenStudy (anonymous):

you're probably facepalming yourself in real life omg

OpenStudy (dtan5457):

i just asked you to subtract x and give me the equation

OpenStudy (dtan5457):

why do you find that confusing lol

OpenStudy (dtan5457):

its as simple as the instructions

OpenStudy (anonymous):

I know ! I know lol like but where do I subtract at? a Calculator? desmo? just take x away like 3y=15? D:

OpenStudy (dtan5457):

ok WATCH KID EXAMPLE X+Z+C+A+D=0 ISOLATE X I HAVE TO REMOVE THE OTHER VARIABLES X=-Z-C-A-D therefore i just SUBTRACT it to the other side. i dont see a need for a calculator

OpenStudy (anonymous):

...

OpenStudy (anonymous):

._. sorry.

OpenStudy (dtan5457):

x-3y=15 -x -x =-3y=15-x

OpenStudy (dtan5457):

can you see that mate

OpenStudy (anonymous):

yes I see it ._.

OpenStudy (dtan5457):

y=1/3x-5

OpenStudy (dtan5457):

how do i isolate y here, please tell me you can tell me something here

OpenStudy (dtan5457):

hint: its a trick question

OpenStudy (dtan5457):

if this takes you more than 1 min then you didn't get my concept lol

OpenStudy (dtan5457):

more than 1 millisecond

OpenStudy (anonymous):

sorry my wifi is like acting stupid

OpenStudy (dtan5457):

its ok but y=1/3x-5 how do i isolate y

OpenStudy (anonymous):

you um you put the y on the other said? =-5-y?